OpenStudy (jenniferjuice):

Mark works at a local fast-food restaurant. He currently makes $350 per week after taxes plus $40 extra cash every fourth week for cleaning the freezer. Mark is trying to save up for a car. He really wants a Mustang. He knows that he can buy a used Mustang in good condition for between $3,600 and $7,200. The equation below represents this situation where x is the number of weeks worked and P is the price of the car: one fourth times forty x plus three hundred fifty x equals P. What is the minimum number of weeks that Mark needs to work to buy his Mustang?

OpenStudy (jenniferjuice):

20 weeks 7 weeks 10 weeks 15 weeks

OpenStudy (anonymous):

1/4(40x) + 350x = p 10x + 350x = 3600 360x = 3600 x = 3600/360 = 10 10 weeks will get him $ 3600